A private plane can make a 125-mile trip in half an hour when flying against a certain headwind. Flying with the same wind, it can make the trip in 25 minutes. Find the speed of the wind and the plane's airspeed.
Answer by checkley77(12844)

D=RT
125=R*.5
R=125/.5
R=250 MPH. AGAINST THE WIND
125=R*25/60
R=125/(25/60)
R=125*60/25
R=4*60=240 MPH
(250-240)/2=10/2=5 MPH IS THE SPEED OF THE WIND.
250-5=245 SPEED OF THE PLANE IN STILL AIR.
240+5=245 DITTO.